HETAV PATEL

Grade 11,

Molality of sulphuric acid solution is 0.2mol/kg find total weight of solution?(in gm)

Molality of sulphuric acid solution is 0.2mol/kg find total weight of solution?(in gm)

Grade:11

3 Answers

Vikas TU
7 years ago
0.2 molality means
0.2 moles of solute are present in 1 kg of solvent.
mass of solution would be=>  mass of solute + mass of solvent.
mass of solute  = molar mass * moles of solute = > 98 * 0.2 = 19.6 gm
Therfore,
mass of soltuion is = 1000+ 19.6 = 1019.6 gm.
Aditya kulkarni
7 years ago
Molality= moles of solute ÷ mass of solventHence moles of solute= 0.2 Mass of solvent =1000gHence, moles of solute *Molecular mass= given massMolar mass for sulphuric acid = 98Hence, 98*0.2= 19.6g (given mass)We know Mass of solution =mass of solute + mass of solvent = 19.6+1000=1019.6g
Aditya Singh chauhan
4 years ago
If the weight of solvent that is 1kg or 1000 g is given in question so we have to calculate total weight of solution.
.2 molal of sulphuric acid
Molecular weight of h2so4 is 98g 
Hence weight of solute is 98g×.2=19.6g
So total weight of solution will be 1000g + 19.6g=1019.6g.ans(cheers)